• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 20:30:55 on February 09, 2011

    Entry number 346242


    keyword=L-HRS: Cosmic Ray Data checkout: issues on TDCs of s1,s2m,cer, ADCs of s2m, V plans of VDC

    As you can see from those graphs bellow, there are several issues:
    1,TDCs of s1,s2m,cer do not show the correct distributions as we have
    last week. What changed since last week is that we have an all-in-one DAQ
    system and the L1A is sent from TS in Bigbite weldment, and people who
    working on FPP re-arranged the ADC&TDC modules in ROC4, where all the
    TDCs of s1,s2m and cer are in. 
    The ADC issues of s2m may or may not relates to the EDTM signal. I asked
    Vince to turn off EDTM signals and take a pure cosmic ray data, after he
    finishes his work.
    I am not sure whether is the wrong timing because of the new DAQ setup,
    or the ROC4 has not been fixed completely. 
    2, V plans of VDC are missing all infomation. The strange thing is that
    even though ROC4 has a problem, it should not affect V plans, since
    ADC&TDC modules of them are in ROC3.

    A copy of this log entry has been emailed to: danez@jlab.org, rom@jlab.org,vasulk@jlab.org, brads@jlab.org

    Figure 1

    Figure 2

    Figure 3

    Figure 4

    Figure 5

    Figure 6

    Figure 7

    Figure 8

    Figure 9

    Figure 10